OpenStudy (anonymous):

If 12 pencils cost 42 cents, how much would 100 pencils cost?

hero (hero):

Find the cost for one pencil, then multiply that by 100

OpenStudy (anonymous):

12/42=x/100

hero (hero):

or you can set up a proportion

OpenStudy (anonymous):

i mean 12/42=100/x

OpenStudy (anonymous):

x=350

OpenStudy (anonymous):

so $3.5

OpenStudy (anonymous):

X=cost of 100 pencils proportion: 12/42=100/x 12(x)=.42(100) 12x=42.00 X=$3.50 cost of 100 pencils

OpenStudy (anonymous):

$3.50

OpenStudy (anonymous):

Thank You, Hero